OpenStudy (anonymous):

Find the volume of a square pyramid with base edges of 48 cm and a slant height of 26 cm. 11,520 cm^3 23,040 cm^3 7,680 cm^3 768 cm^3

OpenStudy (anonymous):

is the 11,520 in the question?

OpenStudy (anonymous):

@vivianrush

OpenStudy (anonymous):

its one of the choices

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I think 7,680 but I'm not sure

OpenStudy (mathstudent55):

 By the Pythagoras theorem, h^2 + 24^2 = 26^2 h^2 = 676 - 576 = 100 h = 10 The volume of a square pyramid is V = (1/3)Ah where A = area of the base and h = height. V = (1/3)(48 cm)^2(10 cm) V = 7,680 cm^3